OpenStudy (anonymous):

Find a * b a = <2, 4> b = <2, 5> Choices: <4,9> 24 <4,20> 16

OpenStudy (anonymous):

you a dot b? 2(2) + 4(5) = 4 + 20 = 24

OpenStudy (anonymous):

If this is a dot b, Sourwing has you covered.

OpenStudy (anonymous):

Thank you, guys!